1. If you had 25.00 mL of an tosic acid solution (HTS) that had a concentration of 4.556 g per liter, how many milliliters
Answer #1

Mass of HTs present = concentration * volume
= 4.556 g/L * 0.0250 L
= 0.1139 g

Mol of HTs = mass / molar mass
= 0.1139 g / 172.2 g/mol
= 6.614*10^-4 mol

The reaction is:
HTs + NaOH -> NaTs + H2O

So,
Mol of NaOH reacting = Mol of HTs
= 6.614*10^-4 mol

Now use:
Mol of NaOH = M(NaOH)*V(NaOH)
6.614*10^-4 mol = 0.396 M * V(NaOH)
V(NaOH) =1.67*10^-3 L
= 1.67 mL

Answer: 1.67 mL